a light beam strikes the surface of glass at a 60° angle to the normal.determine the angle of refraction . given the refracted index of glass and air is 1.5 and 1.0.
1
Expert's answer
2020-12-28T08:35:48-0500


a light beam strikes the surface of glass at a 60° angle to the normal

"\theta_1" = 60o and "\theta_2" - angle of refraction

we use Snell's law for determine angle of refraction - "\frac{sin\theta_1 }{sin\theta_2} = \frac{\nu_1}{\nu_2} = \frac{n_2}{n_2}"

here n - refraction index

n1 = 1.0 for air and n2 = 1.5 for glass


"\frac{sin60^o}{sin\theta_2} = \frac{\frac{3}{2}}{1}" -> "sin\theta_2= \frac{2sin60^o}{3} = \frac{1}{\sqrt{3}}" -> "\theta_2 = 35.3^o"